Solve. A cereal manufacturer uses three large overhead bins to hold the three ingredients in one of its cereal mixes. Bin A delivers a premeasured quantity of dried fruit every 25 minutes, bin B delivers raisins every 15 minutes, and bin C delivers flakes every 10 minutes. If they start the morning shift at the same time, how long beforethey deliver their ingredients at the same time again?
Answer by ankor@dixie-net.com(22740)

A cereal manufacturer uses three large overhead bins to hold the three ingredients in one of its cereal mixes.
Bin A delivers a premeasured quantity of dried fruit every 25 minutes,
bin B delivers raisins every 15 minutes,
and bin C delivers flakes every 10 minutes.
If they start the morning shift at the same time, how long before they deliver
their ingredients at the same time again?
:
Prime factor each
25 = 5 * 5
15 = 5 * 3
10 = 5 * 2
:
factor of LCM = 5 * 5 * 3 * 2 = 150 minutes